OpenStudy (anonymous):

Help!! Giving out medals ;) Find the common ratio of the sequence -164, -82, -41, -20.5,.. A. 2 B. -82 C. 1/2 D. 82

OpenStudy (ahsome):

@gxldmouth, what happened in the first number \(-164\) to become the second number \(-82\)?

OpenStudy (dtan5457):

its evident that the numbers get smaller (technically bigger as it is negatives) right?

OpenStudy (ahsome):

@gxldmouth?

OpenStudy (solomonzelman):

this is how you can find \(\large\color{black}{r }\) (the common ratio): \(\Large\color{blue}{r= a_{n}~\div~a_{n-1}}\)

OpenStudy (danjs):

n1 = -164 n2 = -82 n1 = k * n2 What is the constant k?

OpenStudy (danjs):

rather n2 = k*n1 what is the proportionality constant k?

OpenStudy (solomonzelman):

you can use either of the following: ~ \(\Large\color{blue}{r= a_{2}~\div~a_{1}}\) ~ \(\Large\color{blue}{r= a_{3}~\div~a_{2}}\) ~ \(\Large\color{blue}{r= a_{4}~\div~a_{3}}\)

OpenStudy (solomonzelman):

don't get my notations?

OpenStudy (anonymous):

what do u think the numbers are doing each step hint a. adding two b. subtracting 82 c. dividing by two d. adding 82

OpenStudy (danjs):

same as @SolomonZelman , \(\large\color{blue}{ {\bf a2 ~~\alpha ~~r*a1} }\)

OpenStudy (anonymous):

same as @DanJS and @SolomonZelman but mine is a cheat sheet ;)

OpenStudy (danjs):

I think all the funny 'letters' scared him away

OpenStudy (anonymous):

lol

OpenStudy (solomonzelman):

one more thing I will say: \(\large\color{black}{ a_{_1},~~~~~~~~~~~~~~~~~~~~a_{_2},~~~~~~~~~~~~~~~~~~~~a_{_3},~~~~~~~~~~~~~~~~a_{_4},~~~ }\) \(\large\color{black}{ \Downarrow~~~~~~~~~~~~~~~~~~~~~~~\Downarrow~~~~~~~~~~~~~~~~~~~\Downarrow~~~~~~~~~~~~~~~~~\Downarrow~~~ }\) \(\large\color{black}{ -164,~~~~~~~~~~~~-82,~~~~~~~~~~~~~~~-41,~~~~~~~~~~~~~~-20.5,~~~ }\)

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

i need an answer

OpenStudy (anonymous):

Express the answer in simplest form. -1/3+w=5and 3 over 8 and the answer has to be a mixed fraction

OpenStudy (anonymous):

i will give medal